# Infinite dSprites for Disentangled Continual Learning: Separating Memory Edits from Generalization

### v2

@@ 4.1 Continual learning
Continual learning literature typically focuses on catastrophic forgetting in supervised classification. Parameter isolation methods use dedicated parameters for each task by periodically extending the architecture while freezing already trained parameters (Rusu et al., 2016) or by relying on isolated sub-networks (Fernando et al., 2017). Regularization approaches aim to preserve existing knowledge by limiting the plasticity of the network. Functional regularization methods constrain the network output through knowledge distillation (Li & Hoiem, 2017a) or by using a small set of anchor points to build a functional prior (Pan et al., 2020; Titsias et al., 2020). Weight regularization methods Zenke et al. (2017) directly constrain network parameters according to their estimated importance for previous tasks. In particular, Variational Continual Learning (VCL) by Nguyen et al. (2018) derives the importance estimate by framing continual learning as sequential approximate Bayesian inference. Most methods incorporate regularization into the objective function, but it is also possible to implement it using constrained optimization Lopez-Paz & Ranzato (2017); Aljundi et al. (2019b); Hess et al. (2023); Kao et al. (2021). Finally, replay methods (Rebuffi et al., 2017; Chaudhry et al., 2019; Isele & Cosgun, 2018; Rolnick et al., 2019) retain knowledge through rehearsal. When learning a new task, the network is trained with a mix of new samples from the training stream and previously seen samples drawn from the memory buffer. A specific case of this strategy is generative replay (Shin et al., 2017; Atkinson et al., 2018), where the rehearsal samples are produced by a generative model trained to approximate the data distribution for each class. Finally, continual adaptation approaches rely on a large pre-trained model that is then
@@ 5.1 Regularization methods
In this section, we compare our approach to standard regularization methods: Learning without Forgetting (LwF), Synaptic Intelligence (SI), and Elastic Weight Consolidation (EWC) (Li & Hoiem, 2017b; Zenke et al., 2017; Kirkpatrick et al., 2017). We use implementations from Avalanche, a continual learning library by Lomonaco et al. (2021). We provide details of the hyperparameter choice in the supplementary material. As shown in fig. 4, such regularization methods are ill-equipped to deal with long horizon class-incremental learning scenario and their performance deteriorates rapidly after only 10 tasks.
@@ 5.2 Replay-based methods
Rehearsal can be a viable strategy to maintain high accuracy over hundreds of tasks, but unless the buffer size is bounded, its memory footprint grows rapidly over time. In this section, we investigate the effect of maximum buffer size on performance for standard experience replay with reservoir sampling. While there are replay-based methods that improve on this baseline, we are interested in investigating the fundamental limits of rehearsal over long time horizons and strip away the confounding effects of data augmentation, generative replay, sampling strategies, etc. Figure 4 shows test accuracy for experience replay with different buffer sizes. Storing enough past samples lets the model maintain high test accuracy, but even with a buffer of 20,000 images the performance eventually starts to deteriorate. We provide additional comparisons to Averaged Gradient Episodic Memory (Chaudhry et al., 2018) and Dark Experience Replay (Buzzega et al., 2020) in section A.2.
@@
5.3 Large pre-trained models
With the advent of large vision-language models, parameter-efficient adaptation emerged as a popular approach to continual learning. Typically, it involves pre-training a large model on a large dataset like ImageNet-21k (Russakovsky et al., 2015; Ridnik et al., 2021) and continually adapting a smaller set of parameters in the hope of combining the power of a general-purpose model with the plasticity of adapter methods, such as low rank adaptation or prompt tuning (Hu et al., 2021; Jia et al., 2022). These methods achieve impressive results on natural image benchmarks, such as Split-CIFAR or Split-ImageNet, whose categories largely overlap with the training data of the original foundation model. Nonetheless, in many real-world problems of interest such as remote sensing, astronomy, medical imaging, or specialised vision systems, the data follows completely different distributions. Adapting to domains that are not covered by the pre-training datasets is more difficult and prone to catastrophic forgetting. To demonstrate it, we test Learning to Prompt (L2P), a prototypical continual prompt tuning method introduced by Panos et al. (2023) on our synthetic dataset. As shown in fig. 6 the performance of the 86M parameter ViT-B/16 model on this simple dataset rapidly deteriorates. With this experiment, we demonstrate that despite it synthetic nature, idSprites still presents a considerable challenge for continual learning methods based on pre-trained models.
@@ 5.4 Foundation models
Multimodal foundation models can perform an impressive range of computer vision tasks zero-shot or via in-context learning. Their capabilities include classification, visual question answering, object detection, and optical character recognition. In this section, we test whether GPT-4, a state of the art foundation model introduced by Achiam et al. (2023) is already able to solve the shape re-identification task that lies at the core of idSprites. To this end, we present the model with a query shape in a random position, orientation, and scale. We then show multiple shapes in the canonical from and ask it to pick the exemplar matching the query. Figure 6 shows the average accuracy over 100 multiple-choice questions as a function of the number of possible answers. While the model performs better than chance when there are only a few answers to choose from, we conclude that GPT-4 zero-shot object re-identification capabilities are not yet at a level where it could reliably solve our simple benchmark. The details of the experiment including the exact prompt are provided in section A.3.
@@ 5.5 Do we need equivariance?
By training a network to regress the value of each FoV, we learn a representation that is equivariant to affine transformations. However, we could also take advantage of the shape labels to learn an invariant representation that we could then use to perform classification. In this section, we directly compare equivariant and invariant learning to demonstrate further that learning an equivariant representation is the key to achieving effective continual learning within our framework.
Our baseline for invariant representation learning is based on SimCLR (Chen et al., 2020), a simple and effective contrastive learning algorithm that aims to learn representations invariant to data augmentations. To adapt SimCLR to our problem, we introduce two optimization objectives. The first objective pulls the representation of each training point towards the representation of its exemplar while repelling all other training points. The second objective encourages well-separated exemplar representations by pushing the representations of all exemplars in the current task away from each other. We observed that the first training objective alone is sufficient, but including the second loss term speeds up training. For each task, we train the baseline until convergence. At test time, the class labels are assigned through nearest neighbor lookup in the representation space. Similar to our method, we store a single exemplar per class.
@@
Figure 8: Cumulative test accuracy of Disentangled Continual Learning in the online scenario (every data point is only seen once) and two offline scenarios (3 and 5 training epochs per task).
5.6 Online vs. offline
In all previous experiments, we applied our method in batch mode: we performed multiple training passes over the data for each task. However, efficiently learning from streaming data might require observing each training sample only once to make sure computation is not becoming a bottleneck. This is why we test our method in the online learning regime and compare it to two batch learning scenarios. The results are shown in fig. 8. Unsurprisingly, training for multiple epochs results in better and more robust accuracy on past tasks; it is however worth noting that our method still improves over time in the online learning scenario.
@@ 5.7 One-shot generalization
To evaluate whether the learned regression network can generalize to unseen classes, we perform a one-shot learning experiment. Here, the model is asked to classify transformed versions of shapes it had not previously encountered.
Since the returned class label depends on the exemplars in the buffer, we consider two variants of the experiment, corresponding to generalized and standard one-shot learning. In the first one, we keep the training exemplars in the buffer and add new ones. In the second, the buffer is cleared before including novel exemplars. We also introduce different numbers of test classes. The classification accuracies are presented in Table 1. As expected, keeping the training exemplars in the buffer and adding more test classes makes the task harder. Nevertheless, the accuracy stays remarkably high, showing that the equivariant network has learned a correct and universal mechanism that works even for unseen shapes. This is the essence of our framework.
Table 1: One-shot generalization accuracy of Disentangled Continual Learning on previously unseen shapes after training for 200 tasks (2000 shapes). In the standard variant of the task, the buffer is cleared before adding new exemplars. In the more challenging generalized variant, the exemplars for novel objects are added to the buffer. As expected, keeping the training exemplars in the buffer, as well as including more test classes result in slightly worse classification accuracy.
Test classes
10
100
1000
Standard
1.00
0.98
0.95
Generalized
0.96
0.95
0.93
Figure 10: Some of the incorrectly classified unseen shapes in the open-set recognition task. Top: outputs of the normalization module. Bottom: closest training exemplars.
@@ 6 Discussion
Updating synaptic connections in the human brain upon novel experiences does not interfere with the general knowledge accumulated throughout life. Inspired by this insight, we Disentangled Continual Learning, which splits the continual learning problem into (1) memorizing class-specific information relevant to the task and (2) sequentially training a network that models the general aspects of the problem that apply to all instances. This separation enables disentangled model updating, which allows for explicitly updating class-specific information without destroying information pertinent to other classes and continually learning equivariant representations without catastrophic forgetting. As demonstrated experimentally, such a separation exhibits successful forward and backward transfer and exhibits one-shot generalization and open-set recognition performance.
@@
Reproducibility
We will release idSprites as an open-source Python package and publish the benchmark used in this work. We will also publish all code, training scripts, and configuration files required to reproduce every experiment and generate every figure.
Appendix A Appendix: Experiment Details
@@ A.1 Regularization methods
A.2 Replay-based methods
A.3 Foundation models

### v3

@@ 3.2.2 Discussion
Algorithm 1 Online DCL on idSprites (train)
1:Input: Data $D=\{D_{n}\}_{n=1}^{M},D_{n}=\{(x_{i},y_{i},c_{i})\}_{i=1}^{N_{n}}$
2:Input: Exemplars dictionary $E$
3:Initialize: Memory buffer dictionary $M$
4:Initialize: Canonicalization network $f_{\theta}$
5:for each task $n$ do
6:     for each $(x_{i},y_{i},c_{i})$ in $D_{n}$ do
7:         if $c_{i}$ not in $M$ then
8:              $M[c_{i}]\leftarrow E[c_{i}]$
9:         end if
10:         $T\leftarrow\textsc{ComputeTransformation}(y_{i})$
11:         $\mathcal{L}\leftarrow\textsc{MSE}(f_{\theta}(x_{i}),T)$
[…]
13:     end for
14:end for
Algorithm 2 Online DCL on idSprites (test)
1:Input: Data point to classify $x$
2:Input: Memory buffer dictionary $M$
3:Input: Trained canonicalization network $f_{\theta}$
4:Initialize: Minimum distance $d_{\text{min}}\leftarrow\infty$
5:Initialize: Class label $c\leftarrow\text{None}$
6:Canonicalize $\bar{x}\leftarrow f_{\theta}(x)$
7:for each $c_{i},\hat{x}_{i}$ in $M$ do
8:     $d\leftarrow\textsc{MSE}(\hat{x}_{i},\bar{x})$
9:     if $d<d_{\text{min}}$ then
10:         $d_{\text{min}}\leftarrow d$
11:         $c\leftarrow c_{i}$
12:     end if
13:end for
14:Output: Class label $c$
@@ 4.1 Continual learning
Continual learning literature typically focuses on catastrophic forgetting in supervised classification. Parameter isolation methods use dedicated parameters for each task by periodically extending the architecture while freezing already trained parameters (Rusu et al., 2016) or by relying on isolated sub-networks (Fernando et al., 2017). Regularization approaches aim to preserve existing knowledge by limiting the plasticity of the network. Functional regularization methods constrain the network output through knowledge distillation (Li & Hoiem, 2017a) or by using a small set of anchor points to build a functional prior (Pan et al., 2020; Titsias et al., 2020). Weight regularization methods Zenke et al. (2017) directly constrain network parameters according to their estimated importance for previous tasks. In particular, Variational Continual Learning (VCL) by Nguyen et al. (2018) derives the importance estimate by framing continual learning as sequential approximate Bayesian inference. Most methods incorporate regularization into the objective function, but it is also possible to implement it using constrained optimization Lopez-Paz & Ranzato (2017); Aljundi et al. (2019b); Hess et al. (2023a); Kao et al. (2021). Finally, replay methods (Rebuffi et al., 2017; Chaudhry et al., 2019; Isele & Cosgun, 2018; Rolnick et al., 2019) retain knowledge through rehearsal. When learning a new task, the network is trained with a mix of new samples from the training stream and previously seen samples drawn from the memory buffer. A specific case of this strategy is generative replay (Shin et al., 2017; Atkinson et al., 2018), where the rehearsal samples are produced by a generative model trained to approximate the data distribution for each class. Finally, continual adaptation approaches rely on a large pre-trained model that is then
@@ 5.1 Regularization methods
We compare our approach to standard regularization methods: Learning without Forgetting (LwF), Synaptic Intelligence (SI), and Elastic Weight Consolidation (EWC) (Li & Hoiem, 2017b; Zenke et al., 2017; Kirkpatrick et al., 2017). We use implementations from Avalanche, a continual learning library by Lomonaco et al. (2021). We provide details of the hyperparameter choice in the supplementary material. As shown in fig. 4, such regularization methods are ill-equipped to deal with long horizon class-incremental learning scenario and their performance deteriorates rapidly after only 10 tasks. This finding has been previously described by Lesort et al. (2019) and Lomonaco et al. (2020), and theoretically supported by Knoblauch et al. (2020).
@@ 5.2 Replay-based methods
Rehearsal can be a viable strategy to maintain high accuracy over hundreds of tasks, but unless the buffer size is bounded, its memory footprint grows rapidly over time. In this section, we investigate the effect of maximum buffer size on performance for standard experience replay with reservoir sampling. While there are replay-based methods that improve on this baseline, we are interested in investigating the fundamental limits of rehearsal over long time horizons and strip away the confounding effects of data augmentation, generative replay, sampling strategies, etc. As seen in fig. 4, even with a buffer of 20,000 images, the accuracy of experience replay decreases after only a few dozen tasks. After 500 tasks, the buffer contains merely a few samples per class and the accuracy drops considerably, despite using double the compute resources of alternative methods. This is consistent with previous findings of Wang et al. (2022a), who observe the performance of replay-based methods deteriorates as the buffer size shrinks. Conversely, the accuracy degrades for a set buffer size as the model encounters more tasks. It would be possible to combat forgetting by further expanding the buffer size, but given a limited computational budget, the model would still encounter only a fraction of the buffer during training. This fundamental limitation make replay-based methods impractical when learning over a large number of tasks. For completeness, we provide additional comparisons to Averaged Gradient Episodic Memory (Chaudhry et al., 2018) and Dark Experience Replay (Buzzega et al., 2020) in section B.1.
@@
5.3 Prompt-based methods
With the advent of large vision-language models, parameter-efficient fine-tuning emerged as a popular approach to continual learning. Typically, it involves pre-training a large model on a large dataset like ImageNet-21k (Russakovsky et al., 2015; Ridnik et al., 2021) and continually adapting a smaller set of parameters in the hope of combining the power of a general-purpose model with the plasticity of prompt tuning (Hu et al., 2021; Jia et al., 2022). These methods achieve impressive results on natural image benchmarks, such as Split-CIFAR or Split-ImageNet, whose categories largely overlap with the training data of the original foundation model. Nonetheless, in many real-world problems of interest such as remote sensing, astronomy, medical imaging, or specialised vision systems, the data follows completely different distributions. Adapting to domains that are not covered by the pre-training datasets is more difficult and prone to catastrophic forgetting. To demonstrate it, we test Learning to Prompt (L2P), a prototypical continual prompt tuning method introduced by Panos et al. (2023) on our synthetic dataset. As shown in fig. 6, the performance of the 86M parameter ViT-B/16 model on this simple dataset rapidly deteriorates. With this experiment, we demonstrate that despite it synthetic nature, idSprites still presents a considerable challenge for continual learning methods based on pre-trained models.
@@ 5.4 Foundation models
Multimodal foundation models can perform an impressive range of computer vision tasks zero-shot or via in-context learning. Their capabilities include classification, visual question answering, object detection, and optical character recognition. In this section, we test whether GPT-4, a state of the art foundation model introduced by Achiam et al. (2023) is already able to solve the shape re-identification task that lies at the core of idSprites. To this end, we present the model with a query shape in a random position, orientation, and scale. We then show multiple shapes in the canonical form and ask it to pick the exemplar matching the query. Figure 6 shows the average accuracy over 100 multiple-choice questions as a function of the number of possible answers. While the model performs better than chance when there are only a few answers to choose from, we conclude that GPT-4 zero-shot object re-identification capabilities are not yet at a level where it could reliably solve our simple benchmark. The details of the experiment including the exact prompt are provided in section A.2.
5.5 Disentangled continual learning
The particular implementation of Disentangled Continual Learning for Infinite dSprites is so effective because it separates memorization from generalization by introducing a strong inductive bias, constraining the class of symmetries expected in the data to two-dimensional affine transformations. The crucial advantage of this approach is that continually regressing the parameters of these transformations is not prone to catastrophic forgetting. We speculate this is because the underlying objective remains constant even as new shapes appear. The network needs to learn a general representation and find an effective algorithm for estimating the factors of variation: scale, orientation, and position. Since each shape is observed in numerous combinations of these factors, relying on a shape-specific shortcut would fail, leading the network to learn a more general solution. Furthermore, this solution is refined as new tasks appear, indicating that naive fine-tuning on the FoV regression task in a domain-incremental setting is enough to achieve knowledge accumulation.
We believe this finding can be explained by two main factors. First, gradient-based optimization can overcome catastrophic forgetting and demonstrate knowledge accumulation when data reoccurs over a long sequence, as shown recently by Lesort et al. (2023). While each shape in idSprites is unique, there is perhaps enough similarity within types of shapes to lead to improved FoV regression, thereby boosting classification accuracy. Second, because the underlying task of FoV regression remains constant, it is reasonable to expect the network to exhibit less feature forgetting. According to a study by Hess et al. (2023b), this consistency is likely to contribute to improved knowledge accumulation.
@@ 5.6 Do we need equivariance?
To demonstrate further that learning an equivariant representation is the key to achieving effective continual learning within our framework, we compare equivariant and invariant learning directly. Our baseline for invariant representation learning is based on SimCLR (Chen et al., 2020), a simple and effective contrastive learning algorithm that aims to learn representations invariant to data augmentations. To adapt SimCLR to our problem, we introduce two optimization objectives. The first objective pulls the representation of each training point towards the representation of its exemplar while repelling all other training points. The second objective encourages well-separated exemplar representations by pushing the representations of all exemplars in the current task away from each other. We observed that the first training objective alone is sufficient, but including the second loss term speeds up training. For each task, we train the baseline until convergence. At test time, the class labels are assigned through nearest neighbor lookup in the representation space. Similar to our method, we store a single exemplar per class.
@@ 5.7 One-shot generalization
To evaluate whether the learned regression network can generalize to unseen classes, we perform a one-shot learning experiment. Here, the model is asked to classify transformed versions of shapes it had not previously encountered. Since the returned class label depends on the exemplars in the buffer, we consider two variants of the experiment, corresponding to generalized and standard one-shot learning. In the first one, we keep the training exemplars in the buffer and add new ones. In the second, the buffer is cleared before including novel exemplars. We also introduce different numbers of test classes. The classification accuracies are presented in Table 1. As expected, keeping the training exemplars in the buffer and adding more test classes makes the task harder. Nevertheless, the accuracy stays remarkably high, showing that the equivariant network has learned a correct and universal mechanism that works even for unseen shapes. This is the essence of our framework.
Figure 9: Some of the incorrectly classified unseen shapes in the open-set recognition task. Top row: inputs to the equivariant network. Middle row: outputs of the normalization module. Bottom row: closest training exemplars.
Table 1: One-shot generalization accuracy of Disentangled Continual Learning on unseen shapes after training for 200 tasks (10 shapes per task). In the standard variant, the buffer is cleared before adding novel exemplars. In the more challenging generalized variant, the exemplars for unseen shapes are appended to the buffer. Keeping the training exemplars in the buffer, as well as including more test classes result in slightly worse classification accuracy.
Novel classes added
10 100 1000
Standard 1.00 0.98 0.95
Generalized 0.96 0.95 0.93
@@ 6 Discussion
Updating synaptic connections in the human brain upon novel experiences does not interfere with the general knowledge accumulated throughout life. Inspired by this insight, we introduce Disentangled Continual Learning, which decomposes the continual learning problem into (1) memorizing class-specific information relevant to the task and (2) sequentially training a network that models the general aspects of the problem that apply to all instances. This separation enables explicitly updating class-specific information without destroying information pertinent to other classes and continually learning equivariant representations without catastrophic forgetting. As demonstrated experimentally, a method implementing this spearation exhibits successful forward and backward transfer, one-shot generalization, and open-set recognition.
@@
Acknowledgements
This work was supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. This research utilized compute resources at the Tübingen Machine Learning Cloud, DFG FKZ INST 37/1057-1 FUGG. We thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting SD. This work was supported by the National Centre of Science (Poland) Grants No. 2020/39/B/ST6/01511 and 2022/45/B/ST6/02817.
Appendix A Appendix: Experimental details
@@ A.1 Regularization methods
A.2 Foundation models
@@ A.3 Contrastive baseline
Appendix B Appendix: Additional experiments
B.1 Replay-based methods
B.2 Online vs. offline
In all the main paper experiments, we applied our method in offline mode mode: we performed multiple training passes over the data for each task. However, efficiently learning from streaming data might require observing each training sample only once to make sure computation is not becoming a bottleneck. This is why we test our method in the online learning regime and compare it to two offline learning scenarios. The results are shown in fig. 3. Unsurprisingly, training for multiple epochs results in better and more robust accuracy on past tasks; it is however worth noting that our method still improves over time in the online learning scenario.
Figure 3: Cumulative test accuracy of Disentangled Continual Learning in the online scenario (every image is seen only once) and two offline scenarios (3 and 5 training epochs per task).